# Solving Harder Reinforcement Learning Problems

### v3

@@ A.1 Statement and proof that $n\geq k$ samples are required to unbiasedly estimate $\mathrm{pass@k}$
A sketch of a proof of Theorem 8 can be found in Lehmann and Casella [26].
@@
Let $p=\mathbb{P}\big[[f(x_{i})=1]\big]$ be the probability that any one sample is correct. The $\mathrm{pass@k}$ is the complement of the probability that all of $k$ samples are incorrect, $1-(1-p)^{k}$. Because this is a polynomial in $p$ of order $k$, the result follows immediately from Theorem 8. ∎
@@ A.2 Characterization of the Variance
Our proof of Theorem 1 reveals that the classic Codex formula $\rho(n,c,k)$ is a $U$-statistic (as is our generalization $\rho^{(g)}(n,c,k)$). The concept of $U$-statistics was introduced by Hoeffding, who gave
Theorem 9 (Hoeffding [21]).
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed random variables. Let $h(x_{1},\ldots,x_{k})$ be a symmetric “kernel” with $\mathbb{E}[h(X_{1},\ldots,X_{k})^{2}]<\infty$. Define the parameter $\mu=\mathbb{E}[h(X_{1},\ldots,X_{k})]$ and the $U$-statistic
@@
Further, let $h_{1}(x)=\mathbb{E}[h(x,X_{2},\ldots,X_{k})]$ and define the variance $\zeta_{1}=\text{Var}(h_{1}(X_{1}))$.
If $\zeta_{1}>0$ (the non-degenerate case), then as $n\to\infty$, the $U$-statistic $U_{n}$ is asymptotically normal:
$\sqrt{n}(U_{n}-\mu)\xrightarrow{d}\mathcal{N}(0,k^{2}\zeta_{1}),$ (37)
where $\xrightarrow{d}$ denotes convergence in distribution.
For the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ our kernel is $h(\cdot)\equiv\max(\cdot)$, and we have the following.
@@
As $n\rightarrow\infty$ the variance of $\rho(n,c,k)$ decreases at a rate of $1/n$ with
$\mathbb{E}\left[(\rho(n,c,k)-\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k})^{2}\right]=\frac{k^{2}}{n}\zeta_{1},$ (38)
where $\zeta_{1}$ is a constant defined in Theorem 9 that depends on the model $p(x|\theta)$ and the univariate reinforcement learning reward function $g$.
@@ Appendix C Additional Figures
loo minus one partitioned — a method that uses the same partitioning approach as the previous two, but instead of using the naive estimate (which sets every transformed reward to simple max of the raw reward in a given set of $k$ samples) it uses the $s^{(\mathrm{loo}-1)}$ method applied separately to each disjoint set of $k$ samples, and averages that over all such subsets. In this way, this is a trivial generalization of [39] which extends to $n>k$ by applying the basic method to disjoint subsets and averaging the results. We do not subtract a baseline across sets as this did not improve the variance, possibly because the method within each $k$ already includes a variance reduction baseline.
loo minus one all subsets: $s^{(\mathrm{loo}-1)}$ — our novel estimator of Equation 33 that analytically sums over all subsets of size $k$ of the $n$ samples and uses all appropriate subsets of size k1k-1 to form the variance-reducing baseline that retains unbiasedness, thereby non-trivially generalizing [39] to all $n>k$ with strong variance reduction.
@@
Figure 7: Effect of $k^{\mathrm{opt}}$ (used in our PKPO training) on the rolling $\mathrm{pass@k^{eval}}$ in Gemma2 RL. Setting $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ usually achieves the best $\mathrm{pass@k^{eval}}$. Prior work [39] (which is equivalent to the specific case of $k^{\mathrm{opt}}=n=16$ in our notation) is also shown for comparison, and suffers here presumably due to the larger estimator variance and unreliable gradient (see also Figure 4).

### v4

@@ A.1 Statement and proof that $n\geq k$ samples are required to unbiasedly estimate $\mathrm{pass@k}$
A sketch of a proof of Theorem 8 can be found in Lehmann and Casella [lehmancasella].
@@
It is sufficient to consider a single a fixed and observed correctness function $f$, so that the independence of the $x_{i}$ implies the independence of the correctness events $f(x_{i})=1]$. Let $p=\mathbb{P}\big[[f(x_{i})=1]\big]$ be the probability that any single sample is correct. The $\mathrm{pass@k}$ is defined as the complement of the probability that all $k$ samples are incorrect, which for the specific assumptions adopted in this proof is $1-(1-p)^{k}$. Because this expression is a polynomial in $p$ of degree $k$, the result follows immediately from Theorem 8. ∎
@@ A.2 Characterization of the Variance
Our proof of Theorem 1 identifies the pass@k estimator $\rho(n,c,k)$ as a $U$-statistic. To characterize its variance, we apply Hoeffding’s asymptotic theory.
Theorem 9 (Hoeffding [hoeffding1948]).
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed random variables with distribution $F$. Let $h(x_{1},\ldots,x_{k})$ be a symmetric kernel with $\mathbb{E}[h(X_{1},\ldots,X_{k})^{2}]<\infty$. Define the parameter $\mu=\mathbb{E}_{F}[h(X_{1},\ldots,X_{k})]$ and the $U$-statistic:
@@
Let $h_{1}(x)=\mathbb{E}[h(x,X_{2},\ldots,X_{k})]$ be the projection of the kernel onto a single variable. Hoeffding proved that if $\zeta_{1}=\text{Var}(h_{1}(X_{1}))>0$, then as $n\to\infty$:
$\sqrt{n}(U_{n}-\mu)\xrightarrow{d}\mathcal{N}(0,k^{2}\zeta_{1}).$ (37)
In the standard application of $\mathrm{pass@k}$ we evaluate the estimator on a specific problem defined by a prompt and a correctness oracle. While the true pass rate $\nu$ is unknown to the observer, it is a fixed property of the model-problem pair. Consequently, the correctness outcomes of the generated samples are i.i.d. conditioned on the problem.
The following lemma derives the variance parameter $\zeta_{1}$ under this conditioning. We abuse the notation by allowing the $X_{i}$ to denote correctness.
Lemma 2 (Conditional Variance of the Projection).
Fix a problem instance such that the correctnesses $X_{i}$ are i.i.d. $\text{Bernoulli}(\nu)$. For the $\mathrm{pass@k}$ kernel $h(x_{1},\dots,x_{k})=\max(x_{1},\dots,x_{k})$, the variance of the first-order projection is:
$\zeta_{1}(\nu,k)=\nu(1-\nu)^{2k-1}.$ (38)
Proof.
The projection $h_{1}(x)$ is the expected value of the kernel given the first sample is fixed to $x$, while $X_{2},\dots,X_{k}$ remain random variates drawn from $\text{Bernoulli}(\nu)$.
$h_{1}(x)=\mathbb{E}[\max(x,X_{2},\ldots,X_{k})].$ (39)
We evaluate this for the two possible realizations of $x$:
  1. 1.
    Case $x=1$ (Success): The maximum is 1 regardless of the remaining samples.
    $h_{1}(1)=1.$
  2. 2.
    Case $x=0$ (Failure): The maximum is 0 if and only if all remaining k1k-1 samples fail. Since the remaining samples are i.i.d. with failure probability $(1-\nu)$,
    $h_{1}(0)=1-(1-\nu)^{k-1}.$
The projection $h_{1}(X_{1})$ is thus a binary random variable taking value $h_{1}(1)$ with probability $\nu$ and $h_{1}(0)$ with probability $1-\nu$, so that
$\zeta_{1}$ $=\nu(1-\nu)\left(h_{1}(1)-h_{1}(0)\right)^{2}$
$=\nu(1-\nu)\left(1-[1-(1-\nu)^{k-1}]\right)^{2}$
$=\nu(1-\nu)\left((1-\nu)^{k-1}\right)^{2}$
$=\nu(1-\nu)^{2k-1}.$
We can now substitute this explicit form back into Hoeffding’s general result.
@@
For a fixed problem with pass rate $\nu$, as $n\rightarrow\infty$, the asymptotic variance of the estimator $\rho(n,c,k)$ is:
$\text{Var}(\rho)\approx\frac{1}{n}\left[k^{2}\nu(1-\nu)^{2k-1}\right].$ (40)
@@ Appendix C Additional Figures
loo minus one partitioned — a method that uses the same partitioning approach as the previous two, but instead of using the naive estimate (which sets every transformed reward to simple max of the raw reward in a given set of $k$ samples) it uses the $s^{(\mathrm{loo}-1)}$ method applied separately to each disjoint set of $k$ samples, and averages that over all such subsets. In this way, this is a trivial generalization of [tangloo] which extends to $n>k$ by applying the basic method to disjoint subsets and averaging the results. We do not subtract a baseline across sets as this did not improve the variance, possibly because the method within each $k$ already includes a variance reduction baseline.
loo minus one all subsets: $s^{(\mathrm{loo}-1)}$ — our novel estimator of Equation 33 that analytically sums over all subsets of size $k$ of the $n$ samples and uses all appropriate subsets of size k1k-1 to form the variance-reducing baseline that retains unbiasedness, thereby non-trivially generalizing [tangloo] to all $n>k$ with strong variance reduction.
@@
Figure 7: Effect of $k^{\mathrm{opt}}$ (used in our PKPO training) on the rolling $\mathrm{pass@k^{eval}}$ in Gemma2 RL. Setting $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ usually achieves the best $\mathrm{pass@k^{eval}}$. Prior work [tangloo] (which is equivalent to the specific case of $k^{\mathrm{opt}}=n=16$ in our notation) is also shown for comparison, and suffers here presumably due to the larger estimator variance and unreliable gradient (see also Figure 4).